Jackson had the following information extracted from his records for December 2015.

sales per unit % volume 7000 units

210,000

variable

136500

fixed costs 54000

What is the variable cost per unit?

What is the contribution margin

what is the break even point in units?

What is the margin of safety?

If sales increased by 54000 What is the impact on net income?
Answer #1

SOLUTION

1. Variable cost per unit = Total variable costs / No. of units

= $136,500 / 7,000 units

= $19.50

2. Sales price per unit = $210,000 / 7,000 = $30

Contribution margin = Sales price per unit - Variable cost per unit

= $30 - $19.50 = $10.50

3. Breakeven point in units

= Fixed cost / Contribution margin per unit

= $54,000 / $10.5

= 5,143 units

4. Margin of safety = (Actual sales - Breakeven sales) / Actual sales

=(7,000 - 5,143) / 7,000

= 1,857 / 7,000 = 26.52%

5. Impact on net income =

= Increased sales * Contribution margin ratio

= $54,000 * 35% = $18.900

Net income will increase by $18.900